To study for your upcoming assignment, you have placed two of your favorite physics textbooks, one on top of another, on your table as shown. The mass of book 1 is 600 grams. The mass of book 2 is 300 grams. The coefficient of kinetic friction between book 1 and the table is μ

k(g,1)

=0.2

. The coefficient of static friction between book 1 and book 2 is μ

s(1,2)

=0.3

. Use g=9.81m/s

2

Find, in Newtons, F

max

, the maximum force with with you can push the book at the bottom (book 1) such that both books slide together on the table, without the book on the top (book 2) sliding off


1
Expert's answer
2021-06-30T16:18:48-0400
"(m_1+m_2)a=F-\mu_1(m_1+m_2)g\\
(m_1+m_2)\mu_2 g=F-\mu_1(m_1+m_2)g\\F=(\mu_1+\mu_2)(m_1+m_2)g\\F=(0.2+0.3)(0.6+0.3)9.81=4.41\ N"